OpenStudy (anonymous):

consider the function f(x)=5x^2-17x+6 1. find f(x+h)

OpenStudy (anonymous):

just plug in "x+h" for "x" expand and/or simplify as necessary

OpenStudy (anonymous):

5*(x+h)^2 -17(x+h) +6

OpenStudy (anonymous):

= 5*(x^2+ 2*x*h+ h^2) -17*(x+h) +6 and so on....

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

sure:) that help? was there more to the problem?

OpenStudy (anonymous):

yes for number 2 of the problem it says for h cannot equal 0, find f(x+h)-f(x)/h

OpenStudy (anonymous):

yep, I thought so:)

OpenStudy (anonymous):

so we found f(x+h) now let's add f(x) to it and simplify...

OpenStudy (anonymous):

-f(x) I meant ...

OpenStudy (anonymous):

5*(x^2+ 2*x*h+ h^2) -17*(x+h) +6 - ( 5x^2-17x+6 )

OpenStudy (anonymous):

notice the 5*x^2 's cancel.... what else?

OpenStudy (anonymous):

well just real quick, sorry. how did you get the -f(x) ?

OpenStudy (anonymous):

just made the entirety of f(x) negative... (ie I put a minus sign in front of f(x) )

OpenStudy (anonymous):

- ( 5x^2-17x+6 )

OpenStudy (anonymous):

so -5*x^2 +17*x -6

OpenStudy (anonymous):

oh, okay yest! i see it.

OpenStudy (anonymous):

so simplify the rest and let me know what you get...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i got 2xh+h^2+17h. would i need to further break that down?

OpenStudy (anonymous):

does it just stay in that format?

OpenStudy (anonymous):

well remember, it was all over 'h' (we left that out while we were simplifying the numerator...)

OpenStudy (anonymous):

so there's one more simplification to make... factor the h out of the numerator and h/h =1

OpenStudy (anonymous):

looks like you forgot to multiply a few terms by the '5' too... go back and fix that...

OpenStudy (anonymous):

2xh+h^2-85x-85h+30+17x-6/h can you tell me if so far i'm on the right path?

OpenStudy (anonymous):

you're skipping steps or something

OpenStudy (anonymous):

5*(x^2+ 2*x*h+ h^2) -17*(x+h) +6 - ( 5x^2-17x+6 )

OpenStudy (anonymous):

do the first term on the left 5*(x^2+ 2*x*h+ h^2) = ?

OpenStudy (anonymous):

5x^2+10xh+5h^2 ?

OpenStudy (anonymous):

good:)

OpenStudy (anonymous):

rest is pretty easy

OpenStudy (anonymous):

-17x -17h +6 -5x^2 +17x -6

OpenStudy (anonymous):

17x and 17x cancel 6 and -6 cancel all that's left is -17h and -5x^2

OpenStudy (anonymous):

what you got from the first step : what I got from the 2nd: 5x^2+10xh+5h^2 -17h -5x^2

OpenStudy (anonymous):

Is that the format I would leave it in? I wouldn't need to further break it down?

OpenStudy (anonymous):

combine and you get 5h^2 +10xh -17h

OpenStudy (anonymous):

then we need to put that all over h: (5h^2 +10xh -17h)/h = 5h +10x-17

OpenStudy (anonymous):

Thank you so much. For some reason I was having such a stressful time with this problem and I kept coming up with answers that made no sense.

OpenStudy (anonymous):

:) hope it helped!

OpenStudy (anonymous):

it really did. thank you!

OpenStudy (anonymous):

sure!